# VISO: Robust Underwater Visual-Inertial-Sonar SLAM with Photometric Rendering for Dense 3D Reconstruction

### v1

@@ I Introduction
Underwater simultaneous localisation and mapping (SLAM) is essential for a wide range of tasks, including environmental monitoring, offshore infrastructure inspection, marine archaeology, and autonomous manipulation. SLAM provides underwater vehicles with both accurate pose estimation and reliable environmental perception. However, the underwater environment presents unique characteristics that make accurate localisation and high-fidelity 3D reconstruction challenging. First, sensors such as GPS and Lidars, which are widely used on the ground domain, are unavailable underwater. In addition, visual sensing is severely degraded by light attenuation, scattering, and colour distortion, particularly in turbid waters [rahman2019svin2]. Although multi-beam sonars such as Forward-Looking Sonar (FLS) are unaffected by turbidity, they capture only 2D images, leading to 3D positional ambiguity and pose significant challenges for 3D mapping [hurtos2015fourier, hansen2023using].
To address these challenges, multi-modal sensor fusion strategies have been extensively explored in existing underwater SLAM methods. Cameras, which serve as essential sensors in underwater inspection by providing rich visual information content, have been integrated with other sensing modalities. In particular, visual–inertial systems have been fused with Doppler Velocity Logs (DVLs) [xu2021underwater, xu2025aqua], profiling sonars [rahman2019svin2, rahman2022svin2], and imaging sonars [collado2025opti, pan2025russo], [singh2024opti] to achieve robust underwater localisation and mapping. These solutions can indeed improve the robustness and accuracy of localisation, however, mapping still heavily relies on camera visibility, making 3D scene reconstruction a challenging problem in turbid environments. In contrast, FLS sonars are less susceptible to visually challenging conditions and have been fused with Inertial Measurement Unit (IMU) and DVLs to achieve accurate perception in murky underwater environments [li2018pose], [wang2022virtual], [xu2024diso], [archieri20253dssdf]. However, images captured by FLS suffer from degradation in elevation angle, causing multi-modal SLAM systems to still struggle with full 6-DoF pose estimation and 3D reconstruction.
@@ II RELATED WORK
3D sonar–camera calibration has been rarely explored, [marburg2015extrinsic] is the only work that proposed an extrinsic calibration method between a 3D imaging sonar and an RGB camera using a specific object setup in a laboratory pool, which is complex and time-consuming. 3D sonar shares similar characteristics with LiDAR, and several works on LiDAR–camera calibration have been proposed in [ishikawa2018lidar], [park2020spatiotemporal], [yuan2021pixel]. However, 3D sonar point clouds are considerably more sparse and noisy than LiDAR point clouds, as shown in Fig. 3, and are closer to the 4D radar point clouds. While several works have proposed different 4D radar-camera calibration approaches to obtain the extrinsic parameters [domhof2019extrinsic], [zhuang20254drc], [cao20254d], the significant differences in the measurement range and field of view (FOV) between 3D sonar and 4D radar hinder the direct application of these methods to 3D sonar–camera calibration.
@@ II-B Sonar-based Underwater SLAM
Sonars have been widely used for underwater SLAM. Approaches that rely solely on imaging sonar for underwater SLAM have been proposed in [hurtos2015fourier] and [hansen2023using] to achieve localisation and mosaic generation. To achieve a more robust perception, [li2018pose], [wang2022virtual], [xu2024diso], [archieri20253dssdf] fused imaging sonar, IMU, and DVL to enable robust localisation and mapping. However, imaging sonar data degrade in elevation angle, posing significant challenges for full 6-DoF pose estimation and 3D mapping. To address this issue, [mcconnell2020fusing], [mcconnell2022overhead], [mcconnell2024large] fused two orthogonal imaging sonar datasets to solve elevation ambiguity and achieve large-scale underwater localisation and 3D reconstruction. Nevertheless, these methods face difficulties in orthogonal sonar data association and still struggle to achieve accurate 6-DoF localisation. 3D sonars are a promising sensor for addressing these challenges. Prior works [ferreira20223dupic], [ferreira2025real], and [hansen2005mosaicing] have investigated underwater SLAM using only 3D sonar. However, relying solely on 3D sonar remains fragile due to the inherent sparsity and noise of the data.
@@ II-C Visual-based Underwater SLAM
Cameras play a crucial role in underwater inspection and have been fused with other sensing modalities in visual-based underwater SLAM frameworks. A robust underwater SLAM that fuses a stereo camera, an IMU, and a profiling sonar has been proposed to address visual challenges in underwater localisation [rahman2019svin2], [rahman2022svin2]. Recently, imaging sonar has been used to integrate with a visual–inertial system to mitigate visual degradation [pan2025russo]. While [collado2025opti] proposed fusing segmented camera images with sonar data to enable localisation and 3D reconstruction in highly turbid underwater environments. In addition, cameras have been combined with other sensing modalities such as DVL, barometers, gyroscopes, and pressure sensors to achieve more robust underwater localisation and mapping [xu2021underwater], [xu2025aqua], [song2024turtlmap], [huang2023tightly], [vargas2021robust], [zhao2023tightly]. However, the unique underwater visual challenges hinder the dense 3D reconstruction capability of current visual-based underwater SLAM systems.
@@ III-B Online Extrinsic Calibration
The transformation $\mathbf{T}_{IC}$ from the camera frame to the IMU frame, and the transformation $\mathbf{T}_{CSo}$ from the 3D sonar frame to the camera frame are crucial for tightly-coupled pose estimation. First, $\mathbf{T}_{IC}$ is calibrated using an online calibration method [qin2017vins]. Then $\mathbf{T}_{CSo}$ is calibrated through the following two stages:
@@ III-B1 Coarse Calibration
We first estimate a coarse transformation $\hat{\mathbf{T}}_{CSo}$ using the initial $\mathcal{N}$ camera and 3D sonar poses, where the camera poses are obtained from the visual–inertial system, and the 3D sonar poses are estimated via a coarse 3D sonar odometry, as follows.
@@
$\mathbf{P}_{So_{k}}=\mathbf{T}_{So_{k}So_{i}}\mathbf{P}_{So_{i}}$ (1)
@@
$\mathbf{T}_{So_{i}}=\mathbf{T}_{So_{k}}\mathbf{T}_{So_{k}So_{i}}$ (2)
@@
$\mathbf{T}_{C_{i-1}C_{i}}\mathbf{\hat{T}}_{CSo}=\mathbf{\hat{T}}_{CSo}\mathbf{T}_{So_{i-1}So_{i}}$ (3)
Finally, the coarse extrinsic $\mathbf{\hat{T}}_{CSo}$ is optimized by solving the following nonlinear problem:
@@ III-B2 Refined Calibration
$\mathbf{P}^{w}_{So_{i}}=\mathbf{T}_{WC_{i}}\mathbf{\hat{T}}_{CSo}\mathbf{P}_{So_{i}}$ (5)
@@
$\mathbf{P}^{w}_{C}=\mathbf{T}_{P2P}\mathbf{\hat{P}}^{w}_{C}$ (6)
@@
$\mathbf{T}_{CSo}=\mathbf{T}_{WC_{i}}^{-1}\mathbf{T}_{P2P}\mathbf{T}_{WC_{i}}\mathbf{\hat{T}}_{CSo}$ (7)
@@ III-C1 Compute Surface Features and Normals
Inspired by [adolfsson2022lidar] and [wu2024efear], we first partition the 3D sonar point cloud into $\mathcal{V}$ voxels. For each voxel, we search for its neighbouring points, which include both points inside the voxel and adjacent points from neighbouring voxels. If the number of neighbouring points exceeds a threshold $\gamma$, the voxel is selected as a surface feature, and represented using the mean of its points as $\mathbf{P}_{m}$. Finally, the normal vector $\mathbf{u}_{m}$ of the voxel is calculated as the description of the voxel using principal component analysis (PCA).
@@ III-C2 Scan to Map Tracking
$\mathbf{\hat{P}}_{So_{k}}=\mathbf{\hat{T}}_{So_{k}So_{i}}\mathbf{P}_{So_{i}}$ (8)
@@ III-C4 Feature Point Distance Minimization Residual
$E_{so}=\mathbf{T}_{WI_{k}}\mathbf{T}_{ISo}\mathbf{P}^{k}-\mathbf{T}_{WI_{i}}\mathbf{T}_{ISo}\mathbf{P}^{i}$ (10)
@@ III-D IMU Residual
$\hat{\mathbf{a}}_{t}$ $=\mathbf{a}_{t}+\mathbf{q}^{t}_{bw}\mathbf{g}^{w}+\mathbf{b}_{a_{t}}+\mathbf{n}_{a}$ (11)
$\hat{\boldsymbol{\omega}}_{t}$ $=\boldsymbol{\omega}_{t}+\mathbf{b}_{\omega_{t}}+\mathbf{n}_{\omega}$
@@
$\boldsymbol{\alpha}_{b_{k}b_{k+1}}=\int\int_{t\in[k,{k+1}]}\mathbf{q}^{b_{k}}_{t}(\mathbf{a}_{t}-\mathbf{b}_{a_{t}})dt^{2}$ (12)
$\boldsymbol{\beta}_{b_{k}b_{k+1}}=\int_{t\in[k,{k+1}]}\mathbf{q}^{b_{k}}_{t}(\mathbf{a}_{t}-\mathbf{b}_{a_{t}})dt$
$\mathbf{q}_{b_{k}b_{k+1}}=\int_{t\in[k,{k+1}]}\mathbf{q}^{b_{k}}_{t}\otimes\begin{bmatrix}0\\ \frac{1}{2}\boldsymbol{(\omega}_{t}-\mathbf{b}_{\omega_{t}})\\ \end{bmatrix}dt$
@@
$\mathbf{p}_{wb_{k+1}}$ $=\mathbf{p}_{wb_{k}}+\mathbf{v}_{wb_{k}}\Delta t-\frac{1}{2}\mathbf{g}^{w}\Delta t^{2}+\mathbf{q}_{wb_{k}}\boldsymbol{\alpha}_{b_{k}b_{k+1}}$ (13)
$\mathbf{v}_{wb_{k+1}}$ $=\mathbf{v}_{wb_{k}}-\mathbf{g}^{w}\Delta t+\mathbf{q}_{wb_{k}}\boldsymbol{\beta}_{b_{k}b_{k+1}}$
$\mathbf{q}_{wb_{k+1}}$ $=\mathbf{q}_{wb_{k}}\mathbf{q}_{b_{k}b_{k+1}}$
@@
$\mathcal{R}_{P}$ $=\mathbf{q}_{b_{k}w}(\mathbf{p}_{wb_{k+1}}-\mathbf{p}_{wb_{k}}-\mathbf{v}_{wb_{k}}\Delta t+\frac{1}{2}\mathbf{g}^{w}\Delta t^{2})-\boldsymbol{\alpha}_{b_{k}b_{k+1}}$ (14)
$\mathcal{R}_{V}$ $=\mathbf{q}_{b_{k}w}(\mathbf{v}_{wb_{k+1}}-\mathbf{v}_{wb_{k}}+\mathbf{g}^{w}\Delta t)-\boldsymbol{\beta}_{b_{k}b_{k+1}}$
$\mathcal{R}_{Q}$ $=2[(\mathbf{q}_{b_{k}b_{k+1}})^{-1}\otimes(\hat{\mathbf{q}}_{b_{k}w}\otimes\mathbf{q}_{wb_{k+1}})]_{xyz}$
$\mathcal{R}_{b_{\omega}}$ $=\mathbf{b}_{\omega_{k+1}}-\mathbf{b}_{\omega_{k}}$
$\mathcal{R}_{b_{a}}$ $=\mathbf{b}_{a_{k+1}}-\mathbf{b}_{a_{k}}$
@@
$E_{I}=\begin{bmatrix}\mathcal{R}_{P},\mathcal{R}_{V},\mathcal{R}_{Q},\mathcal{R}_{b_{w}},\mathcal{R}_{b_{a}}\end{bmatrix}^{T}$ (15)
@@ III-E Camera Residual
$\mathbf{P}_{C_{s,j}}=\mathbf{R}_{c_{s}b}\mathbf{R}_{wb_{k}}^{-1}(\mathbf{P}^{w}_{j}-\mathbf{p}_{wb_{k}})+\mathbf{p}_{c_{s}b}$ (17)
@@ III-F Visual-Inertial-Sonar Joint Optimization
$\begin{split}\mathbf{J}(\mathbf{X})=&\sum_{k=1}^{\mathcal{K}}\sum_{j\in k}E_{So}^{j,k^{T}}\mathbf{P}_{So}^{k}E_{So}^{j,k^{T}}+\sum_{k=1}^{\mathcal{K}-1}E_{I}^{k^{T}}\mathbf{P}_{I}^{k}E_{I}^{k}\\ &+\sum_{n=1}^{N=2}\sum_{k=1}^{\mathcal{K}}\sum_{j\in\gamma(n,k)}E_{C}^{n,j,k^{T}}\mathbf{P}_{C}^{k}E_{C}^{n,j,k}\end{split}$ (18)
@@ III-G Sonar Cloud Rendering and Dense Mapping
$\begin{bmatrix}x^{i,j},y^{i,j},z^{i,j},g^{i,j}\end{bmatrix}^{T}=\mathcal{M}^{i,j}=\mathbf{T}_{WI_{i}}\mathbf{T}_{ISo}\mathbf{P}_{So_{i,j}}$ (19)
@@
$g^{i,j}=\mathcal{C}(\pi_{s}(\mathbf{T}_{WI_{i}}\mathbf{T}_{ISo}\mathbf{P}_{So_{i,j}}))$ (20)
@@
Finally, the dense map $\mathcal{M}$ is represented as a mesh using the Truncated Signed Distance Function (TSDF) [vizzo2022vdbfusion].

### v2

@@ I Introduction
Underwater simultaneous localisation and mapping (SLAM) is essential for a wide range of tasks, including environmental monitoring, offshore infrastructure inspection, marine archaeology, and autonomous manipulation. SLAM provides underwater vehicles with both accurate pose estimation and reliable environmental perception. However, the underwater environment presents unique characteristics that make accurate localisation and high-fidelity 3D reconstruction challenging. First, sensors such as GPS and Lidars, which are widely used on the ground domain, are unavailable underwater. In addition, visual sensing is severely degraded by light attenuation, scattering, and colour distortion, particularly in turbid waters [19]. Although multi-beam sonars such as Forward-Looking Sonar (FLS) are unaffected by turbidity, they capture only 2D images, leading to 3D positional ambiguity and pose significant challenges for 3D mapping [10, 8].
To address these challenges, multi-modal sensor fusion strategies have been extensively explored in existing underwater SLAM methods. Cameras, which serve as essential sensors in underwater inspection by providing rich visual information content, have been integrated with other sensing modalities. In particular, visual–inertial systems have been fused with Doppler Velocity Logs (DVLs) [28, 30], profiling sonars [19, 20], and imaging sonars [4, 16], [22] to achieve robust underwater localisation and mapping. These solutions can indeed improve the robustness and accuracy of localisation, however, mapping still heavily relies on camera visibility, making 3D scene reconstruction a challenging problem in turbid environments. In contrast, FLS sonars are less susceptible to visually challenging conditions and have been fused with Inertial Measurement Unit (IMU) and DVLs to achieve accurate perception in murky underwater environments [11], [26], [29], [2]. However, images captured by FLS suffer from degradation in elevation angle, causing multi-modal SLAM systems to still struggle with full 6-DoF pose estimation and 3D reconstruction.
@@ II RELATED WORK
3D sonar–camera calibration has been rarely explored, [12] is the only work that proposed an extrinsic calibration method between a 3D imaging sonar and an RGB camera using a specific object setup in a laboratory pool, which is complex and time-consuming. 3D sonar shares similar characteristics with LiDAR, and several works on LiDAR–camera calibration have been proposed in [17], [31]. However, 3D sonar point clouds are considerably more sparse and noisy than LiDAR point clouds, as shown in Fig. 3, and are closer to the 4D radar point clouds. While several works have proposed different 4D radar-camera calibration approaches to obtain the extrinsic parameters [32], [3], the significant differences in the measurement range and field of view (FOV) between 3D sonar and 4D radar hinder the direct application of these methods to 3D sonar–camera calibration.
@@ II-B Sonar-based Underwater SLAM
Sonars have been widely used for underwater SLAM. Approaches that rely solely on imaging sonar for underwater SLAM have been proposed in [10] and [8] to achieve localisation and mosaic generation. To achieve a more robust perception, [11], [26], [29], [2] fused imaging sonar, IMU, and DVL to enable robust localisation and mapping. However, imaging sonar data degrade in elevation angle, posing significant challenges for full 6-DoF pose estimation and 3D mapping. To address this issue, [13], [14] fused two orthogonal imaging sonar datasets to solve elevation ambiguity and achieve large-scale underwater localisation and 3D reconstruction. Nevertheless, these methods face difficulties in orthogonal sonar data association and still struggle to achieve accurate 6-DoF localisation. 3D sonars are a promising sensor for addressing these challenges. Prior works [5], [6], and [7] have investigated underwater SLAM using only 3D sonar. However, relying solely on 3D sonar remains fragile due to the inherent sparsity and noise of the data.
@@ II-C Visual-based Underwater SLAM
Cameras play a crucial role in underwater inspection and have been fused with other sensing modalities in visual-based underwater SLAM frameworks. A robust underwater SLAM that fuses a stereo camera, an IMU, and a profiling sonar has been proposed to address visual challenges in underwater localisation [19], [20]. Recently, imaging sonar has been used to integrate with a visual–inertial system to mitigate visual degradation [16]. While [4] proposed fusing segmented camera images with sonar data to enable localisation and 3D reconstruction in highly turbid underwater environments. In addition, cameras have been combined with other sensing modalities such as DVL, barometers, gyroscopes, and pressure sensors to achieve more robust underwater localisation and mapping [28], [30], [23], [9], [24]. However, the unique underwater visual challenges hinder the dense 3D reconstruction capability of current visual-based underwater SLAM systems.
@@ III-B Online Extrinsic Calibration
The transformation $\mathbf{T}_{IC}$ from the camera frame to the IMU frame, and the transformation $\mathbf{T}_{CSo}$ from the 3D sonar frame to the camera frame are crucial for tightly-coupled pose estimation. First, $\mathbf{T}_{IC}$ is calibrated using an online calibration method [18]. Then $\mathbf{T}_{CSo}$ is calibrated through the following two stages:
@@ III-B1 Coarse Calibration
We first estimate a coarse transformation $\hat{\mathbf{T}}_{CSo}$ using the initial $n$ camera and 3D sonar poses, where the camera poses are obtained from the visual–inertial system, and the 3D sonar poses are estimated via a coarse 3D sonar odometry, as follows.
@@
$\mathbf{P}_{So_{k}}=\mathbf{T}_{So_{k}So_{i}}\mathbf{P}_{So_{i}}.$ (1)
@@
$\mathbf{T}_{So_{i}}=\mathbf{T}_{So_{k}}\mathbf{T}_{So_{k}So_{i}}.$ (2)
@@
$\mathbf{T}_{C_{i-1}C_{i}}\mathbf{\hat{T}}_{CSo}=\mathbf{\hat{T}}_{CSo}\mathbf{T}_{So_{i-1}So_{i}}.$ (3)
Finally, the coarse extrinsic $\mathbf{\hat{T}}_{CSo}$ is optimized by solving the following problem using nonlinear optimization:
@@ III-B2 Refined Calibration
$\mathbf{P}^{w}_{So_{i}}=\mathbf{T}_{WC_{i}}\mathbf{\hat{T}}_{CSo}\mathbf{P}_{So_{i}}.$ (5)
@@
$\mathbf{P}^{w}_{C}=\mathbf{T}_{P2P}\mathbf{\hat{P}}^{w}_{C}.$ (6)
@@
$\mathbf{T}_{CSo}=\mathbf{T}_{WC_{i}}^{-1}\mathbf{T}_{P2P}\mathbf{T}_{WC_{i}}\mathbf{\hat{T}}_{CSo}.$ (7)
@@ III-C1 Compute Surface Features and Normals
Inspired by [1] and [27], we first partition the 3D sonar point cloud into $\mathcal{V}$ voxels. For each voxel, we search for its neighbouring points, which include both points inside the voxel and adjacent points from neighbouring voxels. If the number of neighbouring points exceeds a threshold $\gamma$, the voxel is selected as a surface feature, and represented using the mean of its points as $\mathbf{P}_{m}$. Finally, the normal vector $\mathbf{u}_{m}$ of the voxel is calculated as the description of the voxel using principal component analysis (PCA).
@@ III-C2 Scan to Map Tracking
$\mathbf{\hat{P}}_{So_{k}}=\mathbf{\hat{T}}_{So_{k}So_{i}}\mathbf{P}_{So_{i}}.$ (8)
@@ III-C4 Feature Point Distance Minimization Residual
$E_{so}=\mathbf{T}_{WI_{k}}\mathbf{T}_{ISo}\mathbf{P}^{k}-\mathbf{T}_{WI_{i}}\mathbf{T}_{ISo}\mathbf{P}^{i},$ (10)
@@ III-D IMU Residual
$\begin{aligned} \hat{\mathbf{a}}_{t}&=\mathbf{a}_{t}+\mathbf{q}^{t}_{bw}\mathbf{g}^{w}+\mathbf{b}_{a_{t}}+\mathbf{n}_{a}\\ \hat{\boldsymbol{\omega}}_{t}&=\boldsymbol{\omega}_{t}+\mathbf{b}_{\omega_{t}}+\mathbf{n}_{\omega}\\ \end{aligned},$ (11)
@@
$\begin{aligned} \boldsymbol{\alpha}_{b_{k}b_{k+1}}=\int\int_{t\in[k,{k+1}]}\mathbf{q}^{b_{k}}_{t}(\mathbf{a}_{t}-\mathbf{b}_{a_{t}})dt^{2}\\ \boldsymbol{\beta}_{b_{k}b_{k+1}}=\int_{t\in[k,{k+1}]}\mathbf{q}^{b_{k}}_{t}(\mathbf{a}_{t}-\mathbf{b}_{a_{t}})dt\\ \mathbf{q}_{b_{k}b_{k+1}}=\int_{t\in[k,{k+1}]}\mathbf{q}^{b_{k}}_{t}\otimes\begin{bmatrix}0\\ \frac{1}{2}\boldsymbol{(\omega}_{t}-\mathbf{b}_{\omega_{t}})\\ \end{bmatrix}dt\end{aligned}.$ (12)
@@
$\begin{aligned} \mathbf{p}_{wb_{k+1}}&=\mathbf{p}_{wb_{k}}+\mathbf{v}_{wb_{k}}\Delta t-\frac{1}{2}\mathbf{g}^{w}\Delta t^{2}+\mathbf{q}_{wb_{k}}\boldsymbol{\alpha}_{b_{k}b_{k+1}}\\ \mathbf{v}_{wb_{k+1}}&=\mathbf{v}_{wb_{k}}-\mathbf{g}^{w}\Delta t+\mathbf{q}_{wb_{k}}\boldsymbol{\beta}_{b_{k}b_{k+1}}\\ \mathbf{q}_{wb_{k+1}}&=\mathbf{q}_{wb_{k}}\mathbf{q}_{b_{k}b_{k+1}}\\ \end{aligned}.$ (13)
@@
[…]
@@
$E_{I}=\begin{bmatrix}\mathcal{R}_{P},\mathcal{R}_{V},\mathcal{R}_{Q},\mathcal{R}_{b_{w}},\mathcal{R}_{b_{a}}\end{bmatrix}^{T}.$ (15)
@@ III-E Camera Residual
$\mathbf{P}_{C_{s,j}}=\mathbf{R}_{c_{s}b}\mathbf{R}_{wb_{k}}^{-1}(\mathbf{P}^{w}_{j}-\mathbf{p}_{wb_{k}})+\mathbf{p}_{c_{s}b},$ (17)
@@ III-F Visual-Inertial-Sonar Joint Optimization
$\begin{split}\mathbf{J}(\mathbf{X})=&\sum_{k=1}^{\mathcal{K}}\sum_{j\in k}E_{So}^{j,k^{T}}\mathbf{P}_{So}^{k}E_{So}^{j,k^{T}}+\sum_{k=1}^{\mathcal{K}-1}E_{I}^{k^{T}}\mathbf{P}_{I}^{k}E_{I}^{k}\\ &+\sum_{n=1}^{N=2}\sum_{k=1}^{\mathcal{K}}\sum_{j\in\gamma(n,k)}E_{C}^{n,j,k^{T}}\mathbf{P}_{C}^{k}E_{C}^{n,j,k}\end{split},$ (18)
@@ III-G Sonar Cloud Rendering and Dense Mapping
$\begin{bmatrix}x^{i,j},y^{i,j},z^{i,j},g^{i,j}\end{bmatrix}^{T}=\mathcal{M}^{i,j}=\mathbf{T}_{WI_{i}}\mathbf{T}_{ISo}\mathbf{P}_{So_{i,j}},$ (19)
@@
$g^{i,j}=\mathcal{C}(\pi_{s}(\mathbf{T}_{WI_{i}}\mathbf{T}_{ISo}\mathbf{P}_{So_{i,j}})),$ (20)
@@
Finally, the dense map $\mathcal{M}$ is represented as a mesh using the Truncated Signed Distance Function (TSDF) [25].